# Geometric theory of quark confinement

### v5

@@ 5 Confinement theorem
under the Virasoro constraint (3.21). These are exactly the conditions of the classical Plateau problem. Therefore, every local minimum of our problem is also the local minimum of the Plateau problem. The difference between solutions could arise in case of self-intersecting loop when there could be several local minima. Our solution selects the one that is additive over the closed parts of the loop (the Goldschmidt solution), whereas the Plateau problem selects the global minimum, which could be topologically different (say, a cylindrical surface bounded by both parts of the loop). According to the theorems [Rado1930, Douglas1931, Gulliver1973], such a disconnected, locally minimal surface always exist. ∎
@@
We must also specify the definition of the perturbative factor $W_{\text{pert}}[C]$ in the Wilson loop. It is uniquely defined as an asymptotic expansion generated by iterating the MM equations [MM1981NPB, Mig83], starting from $W=1$. However, the summation of this asymptotic series is not unique beyond the perturbation expansion, which proceeds in powers of the running coupling constant $\lambda_{\text{eff}}[C]\sim 1/\log(\Lambda_{\text{QCD}}|C|)$. Consequently, non-perturbative terms, such as our dressing factor, appear at double-exponential order:
@@
The ambiguity associated with adding the dressing factor with an arbitrary parameter $\kappa$ corresponds to this inherent ambiguity in perturbative QCD. To be specific, we may define $W_{\text{pert}}[C]$ as the Borel sum of the gluon diagrams in the perturbative expansion. The ambiguity of this Borel sum matches the freedom in the choice of $\kappa$. We must acknowledge that our theory of confinement remains incomplete until a unique definition of $W_{\text{pert}}[C]$ is established, along with the precise relationship between $\kappa$ and the QCD mass scale $\Lambda_{\text{QCD}}$.
@@
Subject to these caveats, we have established a precise correspondence between QCD and the area law. As demonstrated in the following section, this framework yields the standard rising Regge trajectories while avoiding the typical inconsistencies associated with string theory in four dimensions. Our approach relies solely on minimal surfaces—a classical and well-defined subject in modern geometry. While the mathematical theory of minimal surfaces as solutions to the Plateau problem is extensive (briefly summarized in A), it generally does not provide explicit analytic solutions for an arbitrary loop $C$.
@@
Fortunately, the main application of our minimal surface – the rising meson spectrum in QCD– does not require such a general solution. The relevant loops in this case, as we shall see below, simplify, allowing for the analytic solution for the minimal surface and directly leading to linear Regge trajectories..
6 Effective Lagrangian and Regge Trajectories
In this section, we derive the effective Lagrangian for the quark-antiquark pair connected by the confining flux. We describe the confining geometry using the Minkowski Helicoid. It is important to clarify the theoretical justification for this choice. The Helicoid (the ruled surface, made of the rotating stick, Fig. 1) is well-known (see A) as a minimal surface in Euclidean space (where it minimizes area).
Technically, our computation goes along the same lines as that in the classical string theory textbooks [zwiebach2009rotating_open, GreenSchwarzWitten1987]. We repeat these steps here to demonstrate that the dynamics of the string as a physical body do not affect the Regge trajectory in the WKB approximation at large angular momenta, when the masses grow. This WKB method applies to our "holographic string" in the same way as it applies to the physical string considered in these textbooks. The crucial difference is that in our theory we do not have vibrational degrees of freedom to add to the spectrum (and create instabilities in 4D).
The conventional string theory uses the minimal surface as the leading order of the WKB approximation and is forced to add the contribution of the fluctuations of the string shape. In our theory, this is a WKB approximation for the rotation of a pair of quarks under a static force induced by the gluon vacuum.
6.1 The Minkowski Helicoid
The Minkowski Helicoid is parameterized by time $\tau\in[0,T]$ and radius $r\in[-R,R]$, see [docarmo1976]. The time $T$ tends to infinity in the end.
$X(\tau,r)=(\ r\cos\omega\tau,\ r\sin\omega\tau,0,\tau)$ (6.1)
The induced Minkowski metric is $g_{ab}=\text{diag}(1-\omega^{2}r^{2},1)$.
In Minkowski space we are not seeking to minimize the area in the geometric sense, as the area functional is not positive definite and the minimization problem is distinct. Instead, our criterion is algebraic: we require a solution to the Loop Equations for the Wilson loop. This demands the self-duality of the area derivative and the additivity of the area functional. We utilize the fact that the Euclidean Helicoid is a minimal surface for arbitrary angular velocity $\omega$. We obtain the Minkowski Helicoid via a double Wick rotation:
$x_{4}\to it,\qquad\omega_{E}\to-i\omega.$ (6.2)
Under this transformation, the argument of the trigonometric functions in the coordinate parametrization becomes real ($\omega_{E}x_{4}\to\omega t$), mapping the Euclidean Helicoid to a real surface in Minkowski space. Crucially, the condition that $S[C]$ is a zero mode of the Loop Equation depends on algebraic identities (self-duality of the area derivative and additivity). By the Identity Theorem, these analytic relations, established for the Euclidean minimal surface, must hold for its analytic continuation into Minkowski space. Thus, the Minkowski Helicoid remains a valid solution to the QCD Loop Equation, bypassing the ill-posed variational problem of minimizing area in a space with indefinite metric.
Figure 1: The helicoid spanned by rotating $\bar{q}q$ pair connected by a rigid stick (string). This minimal surface bounded by a double helix, was discovered by Meusnier in 1785.
Another important point is that the Helicoid boundary loop (double helix) is infinitely smooth, lacking any cusps or self-intersections. This means that no UV divergent renormalization factors in the perturbative Wilson loop factor $W_{pert}[C]$ could potentially affect our computation of the meson Regge trajectory.
6.2 Semiclassical Motivation: Twisted Boundary Conditions
Before solving the spectral equation, we justify the choice of the Helicoid geometry from the perspective of the semiclassical quark path integral.
Consider the Green’s function for a meson composed of scalar quarks with mass $m$. In the quenched approximation, the effective Minkowski action involves a sum over quark trajectories $C$, weighted by the quark mass term and the area of the confining minimal surface $\Sigma$ bounded by $C$. Neglecting the perturbative factor222the perimeter renormalization terms in that factor we include in the physical quark mass. in the Wilson loop at large loops, relevant to our WKB approximation:
$\mathcal{Z}\sim\int\mathcal{D}C\,\exp\left(-\mathrm{i}m\oint_{C}ds-\mathrm{i}\sigma S_{\mathrm{Mink}}[\Sigma_{C}]\right)$ (6.3)
To extract the spectrum of states with angular momentum $J$, we impose twisted boundary conditions. We require the quark positions at time $t=T$ to be a rotation of their positions at $t=0$ by an angle $\Theta=\omega T$ in the rotation plane $xy$:
$(x+\mathrm{i}y)(T)=e^{\mathrm{i}\omega T}(x+\mathrm{i}y)(0)$ (6.4)
The angular momentum $J$ arises as the thermodynamic conjugate to the twist frequency $\omega$.
The ground state of this system is determined by the saddle point of the effective action. This requires the simultaneous minimization of the boundary length (quark term) and the surface area (gluon term).
  • 1.
    The minimization of the boundary length for a rotating particle yields a helical trajectory in Minkowski spacetime.
  • 2.
    The minimal surface spanning two twisted helical boundaries is the Helicoid.
Thus, the Helicoid is not merely an ansatz; it is the unique geometric solution mandated by the kinematics of angular momentum in the semiclassical limit. The balance between the surface tension $\sigma$ (pulling the quarks inward) and the centrifugal barrier (pushing the twisted trajectories outward) determines the physical radius of the meson.
6.3 Effective Lagrangean
We construct the effective Lagrangian $L$ by combining the bulk contribution from the Helicoid area (coming from our confining factor) with the worldline terms for the massive quarks. The confining factor introduces an effective Minkowski area term. We interpret the phase factor for quark propagation as an effective action, adding the worldline terms $2m\int ds$. The total Lagrangian is:
$L=L_{\text{helicoid}}+L_{\text{quarks}}$ (6.5)
The Helicoid contribution, calculated as the area integral in Minkowski space for a ruled surface with angular velocity $\omega$ and radius $R=v/\omega$, is:
$L_{\text{helicoid}}$ $=-\sigma\int_{-R}^{R}dr\sqrt{1-\omega^{2}r^{2}}$
$=-\frac{\sigma}{\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)$ (6.6)
Here $\sigma$ is the physical string tension, related to the parameter $\kappa$ in the confining factor $\exp\left(-\kappa S[C]\right)$ by the relation $\sigma=2\sqrt{2}\kappa$, which accounts for the normalization of the Hodge-dual area functional. Adding the particle term $2m\sqrt{1-v^{2}}$ as specified for the massive endpoints, the total effective Lagrangian becomes:
$L(v,\omega)=-\frac{\sigma}{\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)-2m\sqrt{1-v^{2}}$ (6.7)
It is worth noting, that this mass term is proportional to the perimeter of the loop in Minkowski space. Therefore, the poterntial perturbative terms proportional to the loop perimeter, are included in this physical quark mass $m$. We are assuming that this physical mass (the bare quark mass plus perturbative corrections) adds up to a finite renormalized mass $m$, which , as the experiment tells us, is much smaller than the QCD mass scale, so that $m^{2}\ll\kappa$.
We proceed to compute the conserved quantities via the Legendre transform. The angular momentum $J$ is the conjugate momentum to the rotation frequency:
$J=\frac{\partial L}{\partial\omega}$ (6.8)
Taking into account that $v$ depends on $\omega$ via the fixed radial extent ($v=\omega R$), the differentiation yields:
$J=\frac{\sigma}{\omega^{2}}(\arcsin v-v\sqrt{1-v^{2}})+\frac{2mv^{2}}{\omega\sqrt{1-v^{2}}}$ (6.9)
The Energy $E$ is defined by the Hamiltonian $E=\omega J-L$. Substituting the expressions for $J$ and $L$ and simplifying the algebra, we obtain:
$E=\frac{2\sigma}{\omega}\arcsin v+\frac{2m}{\sqrt{1-v^{2}}}$ (6.10)
6.4 Balance of quark forces and boundary condition
The relationship between the endpoint velocity $v$ and the string length $R$ is not arbitrary; it is determined by the balance of forces at the string endpoint, which depend on the quark mass $m$. We derive this boundary condition by minimizing the action with respect to the geometric radius $R$ of the loop, keeping the angular velocity $\omega$ fixed.
Recall that the total Lagrangian $L$ is composed of the bulk integral for the helicoid and the boundary terms for the quarks:
$L=-2\sigma\int_{0}^{R}dr\sqrt{1-\omega^{2}r^{2}}-2m\sqrt{1-\omega^{2}R^{2}}$ (6.11)
The equation of motion for the endpoint radius $R$ is given by the stationarity condition $\frac{\partial L}{\partial R}=0$. Differentiating the integral using the Leibniz rule yields the tension force, while differentiating the boundary term yields the centrifugal force:
$\frac{\partial L}{\partial R}=-2\sigma\sqrt{1-\omega^{2}R^{2}}-2m\frac{d}{dR}\left(\sqrt{1-\omega^{2}R^{2}}\right)=0$ (6.12)
Substituting $v=\omega R$ and computing the derivative of the mass term:
@@ 6.4 Balance of quark forces and boundary condition
$-2\sigma\sqrt{1-v^{2}}-2m\left(\frac{-\omega^{2}R}{\sqrt{1-\omega^{2}R^{2}}}\right)=0$ (6.13)
Simplifying the expression, we obtain the exact force balance equation for finite mass:
$\sigma\sqrt{1-v^{2}}=\frac{mv^{2}}{R\sqrt{1-v^{2}}}\implies v^{2}=\frac{\sigma R}{\sigma R+m}$ (6.14)
This equation balances the confining string tension (LHS) against the relativistic centrifugal inertia of the quark (RHS).
We now consider the limit $m\to 0$, and we find the boundary condition
$v^{2}\to 1$ (6.15)
Thus, as the quark mass tends to zero, the endpoint velocity $v$ must approach the speed of light ($v\to 1$).
These equations provide the exact parametric equation for the energy of the classical pair of particles with a given angular momentum and a given radius $R$. To recover the Regge trajectories, we consider the limit of massless quarks ($m\to 0$), where the endpoint velocity approaches the speed of light ($v\to 1$). In this limit, the mass terms vanish, and the system is governed purely by the confining geometry:
$E\approx\frac{2\sigma}{\omega}\frac{\pi}{2},\qquad J\approx\frac{\sigma}{\omega^{2}}\frac{\pi}{2}$ (6.16)
Eliminating $\omega$ between these two relations:
$J=\frac{\sigma(\pi/2)}{(\pi\sigma/E)^{2}}\cdot\frac{\pi}{2}\implies J=\frac{E^{2}}{2\pi\sigma}$ (6.17)
We thus recover the linear Regge trajectory:
$E^{2}=2\pi\sigma J$ (6.18)
This result confirms that the linear confinement arises directly from the kinematics of the Hodge-dual minimal surface in Minkowski space, without invoking independent string vibrations or ad-hoc constants.
@@ 7.1 The Failure of Surface Sums in the Continuum
As the coupling decreases towards the continuum limit, the surface tension decreases, and the “tessellated surface” undergoes a Roughening Transition [Hasenfratz1981, Luscher1981]. Beyond this point, the transverse fluctuations of the surface diverge, and the simple strong coupling expansion breaks down.
More critically, attempts to formulate the Wilson loop as a sum over random surfaces in continuous 4D Euclidean space encounter a fatal geometric instability. The entropy of “crumpled” surface configurations overwhelms the Boltzmann weight of the area action. Consequently, the path integral over surfaces is dominated not by smooth 2D sheets, but by Branched Polymers—tree-like structures with Hausdorff dimension $d_{H}=4$ [Frohlich1985, Durhuus1984]. This “crumpled phase” is unphysical for a string theory of hadrons.
@@ 7.3 Distinct from Nambu-Goto String Models
We do not deny the existence of excited gluonic levels (often termed “hybrid potentials” or “excited flux tubes” in lattice QCD [Juge2003]). However, within our framework, these states are not vibrations of the confining minimal surface itself. The minimal surface is a rigid geometric object fixed by the loop $C$. Excited states correspond to specific operator insertions or gluonic excitations on top of this vacuum background, which would be described by higher-order correlation functions rather than the vacuum expectation value of a single Wilson loop.
@@ 7.4 Decoupling of Confinement and Correlations
7.5 Our 12D Surface vs. AdS/CFT Holography
Our solution shares the fundamental philosophy of the Holographic Principle: the physical observables (Wilson loops) in 4D spacetime are determined by the geometry of a minimal surface in a higher-dimensional "bulk" space. However, the nature of this target space and the mechanism of holography differ fundamentally from the standard AdS/CFT correspondence.
@@
    The Target Space: In AdS/CFT, the holographic dual is typically a curved 5-dimensional anti-de Sitter space (plus a compact manifold), representing a gravitational theory. In our framework, the "bulk" space is a flat 12-dimensional space, $\mathbb{R}^{4}\otimes\mathfrak{su}(2)$, formed by the tensor product of Euclidean spacetime and the canonical $\mathfrak{su}(2)$ subalgebra associated with the self-dual sector (similar to the instanton embedding in arbitrary $N_{c}$).
@@
    Conformal Invariance vs. Confinement: Standard AdS/CFT begins with a conformal field theory; confinement requires breaking this symmetry "by hand" (e.g., via hard/soft walls). Our approach is the inverse: we begin with the Area Law as the fundamental physical input. The theory is inherently confining and non-conformal from the outset, with the mass scale set by the string tension $\sigma$.
Thus, our solution can be viewed as "Gauge Holography" rather than "Gravity Holography." The minimal surface is a hologram of the loop, projected not into a fictitious gravitational dimension, but into the internal color space of the QCD vacuum. This avoids the need to map the glueball spectrum to bulk gravitons, allowing for the clear separation of static confinement and dynamic fluctuations discussed above.
Finally, we observe a striking parallel between the physical principle of color confinement and the geometric principle of holography. Just as color confinement renders the internal color space unobservable—allowing us to interact only with gauge-invariant singlets—the holographic principle renders our 12D target space unobservable, allowing us to interact only with its boundary projection. In this framework, the physical Wilson loop $C$ acts as the unique gauge-invariant observable, representing the projection of the hidden bulk geometry onto the physical world. The extra dimensions are not hidden by compactification, but by the very nature of the measurement process in a confining theory.

### v6

@@ 5 Confinement theorem
under the Virasoro constraint (3.21). These are exactly the conditions of the classical Plateau problem. Therefore, every local minimum of our problem is also the local minimum of the Plateau problem. The difference between solutions could arise in case of self-intersecting loop when there could be several local minima. Our solution selects the one that is additive over the closed parts of the loop (the Goldschmidt solution), whereas the Plateau problem selects the global minimum, which could be topologically different (say, a cylindrical surface bounded by both parts of the loop). According to the theorems [19, 2, 6], such a disconnected, locally minimal surface always exist. ∎
@@
We must also specify the definition of the perturbative factor $W_{\text{pert}}[C]$ in the Wilson loop. It is uniquely defined as an asymptotic expansion generated by iterating the MM equations [12, 14], starting from $W=1$. However, the summation of this asymptotic series is not unique beyond the perturbation expansion, which proceeds in powers of the running coupling constant $\lambda_{\text{eff}}[C]\sim 1/\log(\Lambda_{\text{QCD}}|C|)$. Consequently, non-perturbative terms, such as our dressing factor, appear at double-exponential order:
@@
The ambiguity associated with adding the dressing factor with an arbitrary parameter $\kappa$ corresponds to this inherent ambiguity in perturbative QCD. To be specific, we may define $W_{\text{pert}}[C]$ as the Borel sum of the gluon diagrams in the perturbative expansion. The ambiguity of this Borel sum matches the freedom in the choice of $\kappa$. We must acknowledge that our theory of confinement remains incomplete until a unique definition of $W_{\text{pert}}[C]$ is established, along with the precise relationship between $\kappa$ and the QCD mass scale $\Lambda_{\text{QCD}}$. Such a microscopic definition, leading to explicit solution for the Momentum loop $W[P]$ will be presented in [16].
@@
Subject to these caveats, we have established a precise correspondence between QCD and the area law. As demonstrated in the following section, this framework yields the standard rising Regge trajectories while avoiding the typical inconsistencies associated with string theory in four dimensions. Our approach relies solely on minimal surfaces—a classical and well-defined subject in modern geometry. While the mathematical theory of minimal surfaces as solutions to the Plateau problem is extensive, it generally does not provide explicit analytic solutions for an arbitrary loop $C$.
@@
Fortunately, the main application of our minimal surface – the rising meson spectrum in QCD– does not require such a general solution. The relevant loops in this case, as we shall see in the [16] of this series, simplify, allowing for the analytic solution for the minimal surface and directly leading to linear Regge trajectories.
6 String Scale $\kappa$ and Gluon Condensate
The string mass scale $\kappa$ is not fixed by the MM equation itself. As we have seen, the confining factor $\exp\left(-\kappa S[C]\right)$ passes through the equation without correlation with the perturbative solution $W_{pert}[C]$. This factor reflects the non-perturbative properties of the QCD vacuum rather than the perturbative fluctuations of the gluon field.
The fundamental non-perturbative property of the QCD vacuum is the gluon condensate [21]. We relate the string tension $\kappa$ to this condensate by considering a circular loop $C$ of small radius $r$. According to our solution, $S[C]=2\sqrt{2}\pi r^{2}$. We compute the second area derivative at two opposite points $x_{1}=C(0)$ and $x_{2}=C(\pi)$ on the circle.
At small distances $|x_{12}|=2r$, the Operator Product Expansion (OPE) in asymptotically free QCD is dominated by the leading conformal term in $W_{pert}[C]$:
$\frac{\delta^{2}W_{pert}[C]}{\delta\sigma_{\mu\nu}(1)\delta\sigma_{\lambda\rho}(2)}$ $\propto\langle\mathrm{Tr}F_{\mu\nu}(x_{1})U_{0,\pi}F_{\lambda\rho}(x_{2})U_{\pi,2\pi}\rangle$
$\to\frac{\alpha_{eff}}{4\pi}\frac{I_{\mu\lambda}I_{\nu\rho}-I_{\nu\lambda}I_{\mu\rho}}{|x_{12}|^{4}}W_{pert}[C],$ (6.1)
where $I_{\mu\lambda}=\delta_{\mu\lambda}-2\hat{x}_{\mu}\hat{x}_{\lambda}$ with $\hat{x}=x_{12}/|x_{12}|$. We project this perturbative singularity out by contracting with the tensor $\Pi_{\mu\nu\lambda\rho}=\tfrac{1}{2}(\delta_{\mu\lambda}\delta_{\nu\rho}-\delta_{\nu\lambda}\delta_{\mu\rho})$. Simple tensor algebra confirms that the leading term vanishes: $\mathrm{Tr}(\Pi I\otimes I)=0$.
The remaining term is the gluon condensate contribution.
$\Pi_{\mu\nu\lambda\rho}\mathrm{Tr}F_{\mu\nu}(1)U_{1,2}F_{\lambda\rho}(2)U_{2,1}\to$
$-\frac{g^{2}}{2N_{c}}\langle(G^{a}_{\mu\nu})^{2}\rangle W_{pert}[0].$ (6.2)
This constant contribution must be balanced by the second area derivative of our Hodge-dual factor in the ansatz $W=W_{pert}e^{-\kappa S}$. Using the previously found area derivative (3.51) for a circle , we find:
$\frac{\delta^{2}\exp\left(-\kappa S_{\chi}[C]\right)}{\delta\sigma_{\mu\nu}(0)\delta\sigma_{\lambda\rho}(\pi)}_{\mathrm{circle}}=4\kappa^{2}\eta^{3,\chi}_{\mu\nu}\eta^{3,\chi}_{\lambda\rho}\exp\left(-\kappa S_{\chi}[C]\right)$ (6.3)
@@ 6 String Scale $\kappa$ and Gluon Condensate
$W_{pert}[C]\frac{\delta^{2}\exp\left(-\kappa S[C]\right)}{\delta\sigma_{\mu\nu}(0)\delta\sigma_{\lambda\rho}(\pi)}\to$
$4\kappa^{2}W_{pert}[0]\Pi_{\mu\nu\lambda\rho}\eta^{3,\chi}_{\mu\nu}\eta^{3,\chi}_{\lambda\rho}=16\kappa^{2}W_{pert}[0]$ (6.4)
Comparing these terms, we find (in conventional normalizarion)
$\kappa^{2}=\frac{\pi^{2}}{8N_{c}}\left\langle\frac{\alpha_{s}}{\pi}(G^{a}_{\mu\nu})^{2}\right\rangle$ (6.5)
Our association of $\kappa$ with the gluon condensate supports the physical picture of condensed electric flux originally conjectured by Nambu [17]. One may view the term $\kappa S[C]$ as the action cost of a field configuration with energy density $\langle G^{2}\rangle\sim\Lambda_{QCD}^{4}$ spread over a “pancake” surface of thickness $\Lambda_{QCD}^{-2}$. This estimate interprets the dressing factor $\exp\left(-\kappa S[C]\right)$ as a statistical probability $\exp\left(-\beta\delta E\right)$ for a large quark loop propagating through the QCD vacuum (the dual Meissner effect).
This computation, combined with asymptotic freedom, established the correct scaling of the string tension $\kappa\propto\Lambda_{QCD}^{2}$.
@@ 7.1 The Failure of Surface Sums in the Continuum
As the coupling decreases towards the continuum limit, the surface tension decreases, and the “tessellated surface” undergoes a Roughening Transition [7, 10]. Beyond this point, the transverse fluctuations of the surface diverge, and the simple strong coupling expansion breaks down.
More critically, attempts to formulate the Wilson loop as a sum over random surfaces in continuous 4D Euclidean space encounter a fatal geometric instability. The entropy of “crumpled” surface configurations overwhelms the Boltzmann weight of the area action. Consequently, the path integral over surfaces is dominated not by smooth 2D sheets, but by Branched Polymers—tree-like structures with Hausdorff dimension $d_{H}=4$ [5, 4]. This “crumpled phase” is unphysical for a string theory of hadrons.
@@ 7.3 Distinct from Nambu-Goto String Models
We do not deny the existence of excited gluonic levels (often termed “hybrid potentials” or “excited flux tubes” in lattice QCD [8]). However, within our framework, these states are not vibrations of the confining minimal surface itself. The minimal surface is a rigid geometric object fixed by the loop $C$. Excited states correspond to specific operator insertions or gluonic excitations on top of this vacuum background, which would be described by higher-order correlation functions rather than the vacuum expectation value of a single Wilson loop.
@@ 7.4 Decoupling of Confinement and Correlations
7.5 Our 12D Surface vs. AdS/CFT and Holographic QCD
Our solution shares the fundamental philosophy of the Holographic Principle: the physical observables (Wilson loops) in 4D spacetime are determined by the geometry of a minimal surface in a higher-dimensional "bulk" space. However, the nature of this target space and the mechanism of holography differ fundamentally from the standard AdS/CFT correspondence [13] or Holographic QCD models [23, 20].
@@
    The Target Space: In AdS/CFT and models like Sakai-Sugimoto [20], the holographic dual is a curved 5-dimensional anti-de Sitter space (plus a compact manifold), representing a gravitational theory. In our framework, the "bulk" space is a flat 12-dimensional space, $\mathbb{R}^{4}\otimes\mathfrak{su}(2)$, formed by the tensor product of Euclidean spacetime and the canonical $\mathfrak{su}(2)$ subalgebra associated with the self-dual sector.
@@
    Conformal Invariance vs. Confinement: Standard AdS/CFT begins with a conformal field theory ($\mathcal{N}=4$ SYM). To describe QCD, one must break this symmetry "by hand" via deformations of the metric, such as introducing a hard wall, a soft wall, or compactifying D4-branes on a thermal circle [23]. Our approach is the inverse: we begin with the Area Law as the fundamental physical input. The theory is inherently confining and non-conformal from the outset, with the mass scale set naturally by the string tension $\sigma$ (derived from the gluon condensate in Section 6), without ad hoc geometric deformations.
Thus, our solution can be viewed as "Gauge Holography" rather than "Gravity Holography." The minimal surface is a hologram of the loop, projected not into a fictitious gravitational dimension, but into the internal color space of the QCD vacuum.
7.6 Relation to Scattering Amplitude Duality
Finally, we must address the well-known duality between Wilson loops and scattering amplitudes observed in $\mathcal{N}=4$ Super Yang-Mills theory [1, 3]. In that context, the vacuum expectation value of a polygonal Wilson loop equals the MHV scattering amplitude (squared).
It is crucial to recognize that this duality relies entirely on the dual conformal symmetry of the planar $\mathcal{N}=4$ theory. In a conformal theory, the area law is strictly absent (or divergent), and the Wilson loop is dominated by cusp anomalies and perimeter terms.
In contrast, pure Planar QCD is never conformal. The running coupling constant violates conformal symmetry, leading to dimensional transmutation and the generation of a mass gap (as derived in Section 6). Consequently, the Wilson loop in pure QCD is dominated by the Area Law $\exp\left(-\sigma\text{Area}\right)$, which has no counterpart in the conformal scattering amplitudes of $\mathcal{N}=4$ SYM.
Furthermore, the existing literature on the Wilson loop/amplitude duality typically relies on the AdS/CFT conjecture without solving the underlying equations of motion for the gauge field. In this work, we have solved the MM loop equations [12, 14] directly. Since these are the defining non-perturbative equations of the theory, our solution captures the specific dynamics of asymptotically free QCD, which are physically distinct from the kinematics of conformal Super Yang-Mills.